Use the following information to answer the question: Qs = 4/5 P Qd = 20 - 4/5 P What is the size of consumer surplus if a $2 tax is imposed on each item sold? $72.90 15 items. 90 items. $52.90 $3
Answer #1

When the tax of $2 is imposed on sales, supply curve shifts left and this raises the price consumers pay and reduces the quantity sold

Qs = Qd

(4/5)(P + 2) = 20 - (4/5)P

(4/5)P + (4/5)P + 8/5 = 20

1.6P = 18.4

Price that sellers receive = 11.5

Price that buyers pay = 11.5 + 2 = 13.5

Quantity = 20 - (4/5)*13.5 = 9.2 units

Max price is when Q = 0 so we have P = 20/0.8 = 25

Consumer surplus = 0.5*(max price - current price)*quantity

= 0.5*(25 - 13.5)*9.2

= $52.90

Correct choice is $52.90